# Enhancing Job Salary Prediction with Disentangled Composition Effect Modeling: A Neural Prototyping Approach

### v1

@@ Abstract.
copyright: noneccs: Information systems Data mining
1. Introduction
@@
With the rise of the knowledge economy, the skills that individuals possess play a crucial role in determining job compensation and salary levels (Ng and Feldman, 2014; Dix-Carneiro and Kovak, 2015; Burstein and Vogel, 2017). Understanding the salary-influence of skills is vital for effective recruitment (Sun et al., 2021; Ternikov, 2022; Lovaglio et al., 2018). It helps employers develop competitive salary systems to attract a diverse talent pool. Simultaneously, it guides job seekers align their salary expectations to find suitable jobs. However, despite significant efforts to predict salaries based on job positions (Meng et al., 2018, 2022a) and talent demographics (Lazar, 2004), there remains a shortage of effective methods to accurately model salary-influencing patterns based on skill compositions.
@@
Indeed, the market’s numerous skills create a vast combination space, with different jobs requiring unique skill sets. Analyzing their salary-influencing patterns, given the supply-demand relationship among various roles, is a complex set modeling problem. In recent years, neural networks have advanced set-based modeling by capturing intricate set-outcome relationships. Existing progress, represented by DeepSets (Zaheer et al., 2017) and Set Transformer (Lee et al., 2019), typically learn element embeddings and employ set-pooling layers to extract permutation-invariant set representations in an end-to-end way. This brings high expressiveness in capturing element interactions and remarkable prediction accuracy. However, the lack of intrinsic explainability limits their ability to offer clear insights into the effects of skill compositions.
@@
As a data structure with unique characteristics and internal collaboration mechanisms, sets pose unique challenges to model explanations. An illustrative example is provided in Figure 1. Specifically, (A) The collective effect of set elements, rather than individual impacts, significantly influences outcomes. For instance, as AI developers, talents possessing a comprehensive skill package tend to get higher salaries. Since a single skill cannot solely justify the compensation, common explanation paradigms (Sundararajan et al., 2017; Sun et al., 2021) focusing on individual input impacts struggle to provide meaningful insights into salary determination. Symbolizing and quantifying the collective semantic influences within a set presents a notable challenge. (B) A set can be a combination of multi-faceted semantics from different interactions among skills. For instance, a front-end development position in an AI company may also require some AI knowledge, leading to a higher salary than a typical front-end role but not as high as an AI specialist. Intuitively, the salary is jointly determined by the reflected extent and the overall market value of these global semantics. However, disentangling the global semantics from various mixtures and modeling their interaction to influence the entire set’s outcome is a challenging task. (C) Unlike structured grid data like images, sets are discrete, conceptual, and unordered. These high-level features arise from semantic interactions among elements, rather than spatial arrangements. Thus, existing methods (Nauta et al., 2021; Chen et al., 2019; Li et al., 2018), which rely on continuous representations for spatial visualization, face limitations with sets. Providing quantitative explanations for unique set data remains unsolved.
Figure 1. An illustrative example on skill-salary influence.
@@
To tackle these challenges, we propose a novel, intrinsically explainable set-based neural prototyping approach, namely Local-Global Subset DisEntangling Set Network, called LGDESetNet, which models outcomes based on learning multi-faceted input subsets and influential prototypical sets. With LGDESetNet, we realize explainable and accurate salary prediction while disentangling salary-influential skill sets from both an instance-based local view and a task-based global view. In particular, to learn symbolic set influence, LGDESetNet models discrete and permutation-invariant set-formed patterns in a differentiable end-to-end manner. Specifically, we propose a disentangled discrete subset selection layer to identify local subsets reflecting different semantics. In particular, an element co-occurrence graph density regularization is designed for enhancing dense and disentangled extraction. Next, we propose a set-oriented discrete prototype learning method to extract globally influential sets, which generates outputs in a transparent way based on the semantic interactions between the extracted subsets and global prototypes. To flexibly process external job contexts (e.g., time, city), we design two optional fusion layers to integrate skill-wise and set-wise side information in an explainable way. Finally, we conduct extensive experiments on four real-world job salary datasets. The experimental results clearly show that LGDESetNet outperforms state-of-the-art self-explainable models and salary prediction models. Quantitative and qualitative analyses underscore our method’s superior explainability ability to identify influential skill sets in the labor market and quantify their impact on salaries.
@@ 2. Related Work
This paper’s related work falls into three categories: data-driven salary prediction, set modeling, and neural network explainability. We will discuss each category in detail and highlight the differences between our work and existing studies.
2.1. Data-driven Salary Prediction
@@
In recent years, growing efforts have been made to explore salary-influential factors from a data-driven perspective (Quan and Raheem, 2022). Despite examining various factors (Autor and Dorn, 2013; Dix-Carneiro and Kovak, 2015; Burstein and Vogel, 2017; Sun et al., 2025), quantitative salary prediction remains challenging. The evolution of machine learning offers promising solutions for dynamic and effective salary prediction. For example, Lazar et al. (Lazar, 2004) adopted the classical Support Vector Machine (SVM) on talent demographics data to predict incomes. Meng et al. (Meng et al., 2018) proposed a holistic matrix factorization approach for salary benchmarking based on company and job position information. Meng et al. (Meng et al., 2022a) further designed a nonparametric Dirichlet process to capture the latent skill distributions. In (Sun et al., 2021), a cooperative composition deep model is introduced for salary prediction by market-aware skill valuation. Fang et al. (Fang et al., 2023) developed a pre-training method for recruitment tasks, achieving BERT-comparable performance in the downstream salary prediction task. Unlike existing works, our study focuses on the skill composition effect on job salary and proposes a self-explainable deep set model to provide explainable insights into skill-based salary-influencing patterns. Moreover, we holistically consider both local and global job contexts’ influences (e.g, skill proficiency, base city) on skill-salary prediction.
2.2. Set Modeling
@@
In data mining, frequent pattern mining (Li et al., 2022) is used to discover frequently appearing sets within datasets. Neural network-based methods face challenges with set-structured inputs due to their unordered nature. DeepSets (Zaheer et al., 2017) is the prominent work, which proposed a unified framework based on the set-pooling method to compress any unordered set into a single embedding vector. Based on this framework, (Murphy et al., 2018) designed a learnable Janossy pooling layer to enhance the model flexibility. Another research line investigated the attention mechanism to discover the interactions between set elements (Lee et al., 2019; Maron et al., 2020; Hirsch and Gilad-Bachrach, 2021). SetNorm (Zhang et al., 2022b) further proposed a novel set normalization to increase model depths. Some domain-specific studies can be also viewed as special cases of learning sets, including point cloud (Qi et al., 2017; Xu et al., 2018; Wu et al., 2024; Chen et al., 2024), graphs (Maron et al., 2018; Kim et al., 2022), hypergraph (Chien et al., 2021) and images (Yao et al., 2024; Gordon et al., 2019). However, few studies have focused on explaining the set-based modeling process. The closest work to ours is Set-Tree (Hirsch and Gilad-Bachrach, 2021), which provides instance-wise explanations by measuring the frequency each item occurs in the attention-sets. In contrast, our work provides both multi-faceted local explanations and global prototypical examples with explicit quantitative importance weights.
2.3. Neural Network Explainability
@@
Neural network explainability aims to uncover important local or global data patterns for model predictions (He et al., 2024; Li et al., 2024; Guo et al., 2025; Li et al., 2025). One research line is post-hoc explainability, which employs additional models like gradient-based (Selvaraju et al., 2017; Ancona et al., 2017) and perturbation-based (Lundberg and Lee, 2017) methods to explain black-box models. However, such post-hoc methods fail to provide intrinsic model insights in many cases (Rudin, 2019; Schröder et al., 2023; Sundararajan et al., 2017). Another approach is building self-explainable models, leveraging regularization (Lemhadri et al., 2021; Li et al., 2024) or attention mechanisms (Miao et al., 2022) to identify critical input patterns. Moreover, some explored prototype learning (Chen et al., 2019) or case-based reasoning (Kolodner, 1992), mimicking human problem-solving by matching representative examples (Van Gog and Rummel, 2010). PrototypeDNN (Li et al., 2018) and ProtoPNet (Chen et al., 2019)are pioneering works in prototype-based image classification. Subsequent work (Ma et al., 2024; Wang et al., 2021; Rymarczyk et al., 2021; Pach et al., 2025) aims to make explanations more concise and understandable. Prototype learning has been applied to sequences (Ming et al., 2019; Hong et al., 2020; Rajagopal et al., 2021), graphs (Zhang et al., 2022a; Seo et al., 2024; Dai and Wang, 2025), imitation learning (Jiang et al., 2023), traffic flow (Fauvel et al., 2023), and reinforcement learning (Sun et al., 2025; Kenny et al., 2022). As aforementioned, explaining the set-based modeling process remains underexplored. This work focuses on processing set-structured data by disentangling and prototyping relationships among set entities.
Table 1. Major notations in this paper.
@@
$C$ The set-wise external information (job context).
@@
$\gamma^{\text{s}}_{i}$ The multi-hot selection vector of the prototype $P_{i}$.
$\gamma^{\text{lv}}_{i}$ The skill-wise weights of the prototype $P_{i}$.
$\gamma^{\text{sal}}_{i}$ The salary effect weight of the prototype $P_{i}$.
$E^{\text{p}}_{i}$ The continuous embeddings of the prototype $P_{i}$.
@@
$Z^{\text{s}}_{h}$ The projected embeddings of the subset $S_{h}$.
$Z^{\text{p}}_{i}$ The projected embeddings of the prototype $P_{i}$.
3. Problem Formulation
We utilize real-world job posting data (Sun et al., 2023; Meng et al., 2022b). Each job posting consists of the job salary range and descriptive external information, such as company, city, time, and skill level requirements. To mine the interaction among skill compositions, we formulate each job posting as a tuple $(X,C,Y)$, where $X$ represents a skill set, $C$ represents set-wise external information (e.g., job contexts such as time, city), and $Y$ represents its salary outcome. In particular, $X$ is an unordered set of skills $X=\{x_{1},x_{2},...,x_{n}\}$, where $n$ indicates the number of skills. Each skill $x_{i}$ is drawn from a universal set $U$. $N_{total}=|U|$ indicates the total number of skills across the dataset. In many real-world scenarios, each skill $x_{i}$ is possibly associated with skill-wise external information $lv_{i}$ (e.g., proficiency level).
Formally, we define the task of this paper as a skill-based explainable salary prediction problem. Using training data, it aims to train a neural network model $f$ that can predict salary outcomes based on skill sets with associated external information and provides explainable insights into the reasoning process. Additionally, $f$ should be permutation invariant for set processing, implying that $f(X)=f(\pi(X))$ for any permutation $\pi$ of the indices ${1,2,\ldots,|X|}$ and all $X$ in the input domain. Table 1 summarizes the major notations used in this paper.
4. Methodology
@@
Figure 2. The architecture of LGDESetNet.
4.1. Framework Overview
@@
4.2. Disentangled Discrete Subset Selection
@@
4.2.1. Multi-view Subset Selection Network
@@
First, LGDESetNet uses a permutation-invariant encoder to embed each set’s overall semantic features. The input skill vector $V_{X}\in\mathbb{R}^{n}$ can be in any permutation order of inner skills, where $n$ denotes the size of $X$. The representation $\phi(x_{i})$ for each skill in $U$ is learned as an embedding table $E_{total}\in\mathbb{R}^{N_{total}\times d}$, where $d$ represents the embedding dimension. Next, a permutation-invariant pooling function $\text{pool}(\cdot)$ aggregates the input set’s skill representations into the set representation. We introduce the attention mechanism to capture the interactions of skills to different semantics, the skill and set’s representations are projected into $H$ distinct subspaces. For the $h$-th view, each row in $V_{h}^{\text{K}},V_{h}^{\text{V}}\in\mathbb{R}^{n\times d_{h}}$ denotes the key and value vectors of the corresponding skill in $V_{X}$, and $V_{h}^{\text{Q}}\in\mathbb{R}^{d_{h}}$ denotes the query projected from the set embedding. Using the dot-product attention mechanism (Vaswani et al., 2017), we calculate the similarity of each skill in each view as:
(1) $V^{\text{E}}_{h}=V_{h}^{\text{Q}}(V_{h}^{\text{K}})^{T}/\sqrt{d_{h}}.$
Each entry $V^{\text{E}}_{h}\in\mathbb{R}^{N}$ models a skill’s interaction with the $h$-th view.
While activation functions can be used to project interaction weights to the range [0, 1], the continuous interaction score hinders the network’s ability to model in an explicit binary manner. To address this issue, we introduce Gumbel-Sigmoid (Gumbel, 1948; Geng et al., 2020) to enable differentiable skill selection. Specifically, we incorporate Gumbel noise with the Sigmoid function and output the selected skills for each view’s subset as:
(2) $\begin{split}A^{\text{s}}_{h}=\frac{\exp((V^{\text{E}}_{h}+G_{0})/\tau)}{\exp((V^{\text{E}}_{h}+G_{0})/\tau)+\exp((G_{1})/\tau)},\\ u_{i}\sim\text{Uniform}(0,1),i\in\{0,1\},\\ G_{i}=-\log(-\log(u_{i})),i\in\{0,1\},\end{split}$
where $G_{0}$ and $G_{1}$ are two i.i.d Gumbel noises, the temperature $\tau$ controls the smoothness of the sampling function. $A^{\text{s}}_{h}$ becomes a multi-hot vector for selecting a subset $S_{h}$ when $\tau$ approaches 00 and becomes a continuous vector as $\tau\rightarrow\infty$. In the training loop, we gradually reduce $\tau$ to learn the multi-hot representation.
4.2.2. Skill-wise External Information Fusion

### v3

@@ Abstract
[…]
1 Introduction
@@
With the rise of the knowledge economy, the skills that individuals possess play a crucial role in determining job compensation and salary levels [1, 2, 3]. Understanding the salary-influence of skills is vital for effective recruitment [4, 5, 6]. It helps employers develop competitive salary systems to attract a diverse talent pool. Simultaneously, it guides job seekers align their salary expectations to find suitable jobs. However, despite significant efforts to predict salaries based on job positions [7, 8] and talent demographics [9], there remains a shortage of effective methods to accurately model salary-influencing patterns based on skill compositions.
@@
Indeed, the market’s numerous skills create a vast combination space, with different jobs requiring unique skill sets. Analyzing their salary-influencing patterns, given the supply-demand relationship among various roles, is a complex set modeling problem. In recent years, neural networks have advanced set-based modeling by capturing intricate set-outcome relationships. Existing progress, represented by DeepSets [10] and Set Transformer [11], typically learn element embeddings and employ set-pooling layers to extract permutation-invariant set representations in an end-to-end way. This brings high expressiveness in capturing element interactions and remarkable prediction accuracy. However, the lack of intrinsic explainability limits their ability to offer clear insights into the effects of skill compositions.
@@
As a data structure with unique characteristics and internal collaboration mechanisms, sets pose unique challenges to model explanations. An illustrative example is provided in Figure 1. Specifically, (A) The collective effect of set elements, rather than individual impacts, significantly influences outcomes. For instance, as AI developers, talents possessing a comprehensive skill package tend to get higher salaries. Since a single skill cannot solely justify the compensation, common explanation paradigms [12, 4] focusing on individual input impacts struggle to provide meaningful insights into salary determination. Symbolizing and quantifying the collective semantic influences within a set presents a notable challenge. (B) A set can be a combination of multi-faceted semantics from different interactions among skills. For instance, a front-end development position in an AI company may also require some AI knowledge, leading to a higher salary than a typical front-end role but not as high as an AI specialist. Intuitively, the salary is jointly determined by the reflected extent and the overall market value of these global semantics. However, disentangling the global semantics from various mixtures and modeling their interaction to influence the entire set’s outcome is a challenging task. (C) Unlike structured grid data like images, sets are discrete, conceptual, and unordered. These high-level features arise from semantic interactions among elements, rather than spatial arrangements. Thus, existing methods [13, 14, 15], which rely on continuous representations for spatial visualization, face limitations with sets. Providing quantitative explanations for unique set data remains unsolved.
Figure 1: An illustrative example on skill-salary influence.
@@
To tackle these challenges, we propose a novel, intrinsically explainable set-based neural prototyping approach, namely Local-Global Subset DisEntangling Set Network, called LGDESetNet, which models outcomes based on learning multi-faceted input subsets and influential prototypical sets. With LGDESetNet, we realize explainable and accurate salary prediction while disentangling salary-influential skill sets from both an instance-based local view and a task-based global view. In particular, to learn symbolic set influence, LGDESetNet models discrete and permutation-invariant set-formed patterns in a differentiable end-to-end manner. Specifically, we propose a disentangled discrete subset selection layer to identify local subsets reflecting different semantics. In particular, an element co-occurrence graph density regularization is designed for enhancing dense and disentangled extraction. Next, we propose a set-oriented discrete prototype learning method to extract globally influential sets, which generates outputs in a transparent way based on the semantic interactions between the extracted subsets and global prototypes. To flexibly process external job contexts (e.g., time, city), we design two optional fusion layers to integrate skill-wise and set-wise side information in an explainable way. Finally, we conduct extensive experiments on four real-world job salary datasets. The experimental results clearly show that LGDESetNet outperforms state-of-the-art self-explainable models and salary prediction models. Quantitative and qualitative analyses underscore our method’s superior explainability ability to identify influential skill sets in the labor market and quantify their impacts on salary.
@@ 2 Related Work
This paper’s related work falls into three categories: data-driven salary prediction, set modeling, and neural network explainability. We will discuss each in detail and highlight the differences between our work and existing studies.
2.1 Data-driven Salary Prediction
@@
In recent years, growing efforts have been made to explore salary-influential factors from a data-driven perspective [16]. Despite examining various factors [17, 2, 3, 18], quantitative salary prediction remains challenging. The evolution of machine learning offers promising solutions for dynamic and effective salary prediction. For example, Lazar et al. [9] adopted the classical Support Vector Machine (SVM) on talent demographics data to predict incomes. Meng et al. [7] proposed a holistic matrix factorization approach for salary benchmarking based on company and job position information. Meng et al. [8] further designed a nonparametric Dirichlet process to capture the latent skill distributions. In [4], a cooperative composition deep model is introduced for salary prediction by market-aware skill valuation. Fang et al. [19] developed a pre-training method for recruitment tasks, achieving BERT-comparable performance in the downstream salary prediction task. Unlike existing works, our study focuses on the skill composition effect on job salary and proposes a self-explainable deep set model to provide explainable insights into skill-based salary-influencing patterns. Moreover, we holistically consider both local and global job contexts’ influences (e.g, skill proficiency, base city) on skill-salary prediction.
2.2 Set Modeling
@@
In data mining, frequent pattern mining [20] is used to discover frequently appearing sets within datasets. Neural network-based methods face challenges with set-structured inputs due to their unordered nature. DeepSets [10] is the prominent work, which proposed a unified framework based on the set-pooling method to compress any unordered set into a single embedding vector. Based on this framework, [21] designed a learnable Janossy pooling layer to enhance the model flexibility. Another research line investigated the attention mechanism to discover the interactions between set elements [11, 22, 23]. SetNorm [24] further proposed a novel set normalization to increase model depths. Some domain-specific studies can be also viewed as special cases of learning sets, including point cloud [25, 26, 27, 28], graphs [29, 30], hypergraph [31] and images [32, 33]. However, few studies have focused on explaining the set-based modeling process. The closest work to ours is Set-Tree [23], which provides instance-wise explanations by measuring the frequency each item occurs in the attention-sets. In contrast, our work provides both multi-faceted local explanations and global prototypical examples with explicit quantitative importance weights.
2.3 Neural Network Explainability
@@
Neural network explainability aims to uncover important local or global data patterns for model predictions. One research line is post-hoc explainability, which employs additional models like gradient-based [34, 35] and perturbation-based [36] methods to explain black-box models. However, such post-hoc methods fail to provide intrinsic model insights in many cases [37, 38, 12]. Another approach is building self-explainable models, leveraging regularization [39] or attention mechanisms [40] to identify critical input patterns. Moreover, some explored prototype learning [14] or case-based reasoning [41], mimicking human problem-solving by matching representative examples [42]. PrototypeDNN [15] and ProtoPNet [14]are pioneering works in prototype-based image classification. Subsequent work [43, 44, 45, 46] aims to make explanations more concise and understandable. Prototype learning has been applied to sequences [47, 48, 49], graphs [50, 51], imitation learning [52], traffic flow [53], and reinforcement learning [18, 54]. As aforementioned, explaining the set-based modeling process remains underexplored. This work focuses on processing set-structured data by disentangling and prototyping relationships among set entities.
Table 1: Major notations in this paper.
@@
$C$ The set-wise external information.
@@
$\gamma^{\text{s}}_{i}$ The multi-hot selection vector of $P_{i}$.
$\gamma^{\text{lv}}_{i}$ The skill-wise weights of $P_{i}$.
$\gamma^{\text{sal}}_{i}$ The salary effect weight of $P_{i}$.
$E^{\text{p}}_{i}$ The continuous embeddings of $P_{i}$.
@@
$Z^{\text{s}}_{h}$ The projected embeddings of $S_{h}$.
$Z^{\text{p}}_{i}$ The projected embeddings of $P_{i}$.
3 Problem Formulation
We utilize real-world job posting data [55, 56]. Each job posting consists of the job salary range and descriptive external information, such as company, city, time, and skill level requirements. To mine the interaction among skill compositions, we formulate each job posting as a tuple $(X,C,Y)$, where $X$ represents a skill set, $C$ represents set-wise external information (e.g., job contexts such as time, city), and $Y$ represents its salary outcome. In particular, $X$ is an unordered set of skills $X=\{x_{1},x_{2},...,x_{n}\}$, where $n$ indicates the number of skills. Each skill $x_{i}$ is drawn from a universal set $U$. $N_{total}=|U|$ indicates the total number of skills across the dataset. In many real-world scenarios, each skill $x_{i}$ is possibly associated with skill-wise external information $lv_{i}$ (e.g., proficiency level).
Formally, we define the task of this paper as a skill-based explainable salary prediction problem. Using training data, it aims to train a neural network model $f$ that can predict salary outcomes based on skill sets with associated external information and provides explainable insights into the reasoning process. Additionally, $f$ should be permutation invariant for set processing, implying that $f(X)=f(\pi(X))$ for any permutation $\pi$ of the indices ${1,2,\ldots,|X|}$ and all $X$ in the input domain. Table 1 summarizes the major notations.
4 Methodology
@@
Figure 2: The architecture of LGDESetNet.
4.1 Framework Overview
@@
4.2 Disentangled Discrete Subset Selection
@@
4.2.1 Multi-view Subset Selection Network
@@
First, LGDESetNet uses a permutation-invariant encoder to embed each set’s overall semantic features. The input skill vector $V_{X}\in\mathbb{R}^{n}$ can be in any permutation order of inner skills, where $n$ denotes the size of $X$. The representation $\phi(x_{i})$ for each skill in $U$ is learned as an embedding table $E_{total}\in\mathbb{R}^{N_{total}\times d}$, where $d$ represents the embedding dimension. Next, a permutation-invariant pooling function $\text{pool}(\cdot)$ aggregates the input set’s skill representations into the set representation. We introduce the attention mechanism to capture the interactions of skills to different semantics, the skill and set’s representations are projected into $H$ distinct subspaces. For the $h$-th view, each row in $V_{h}^{\text{K}},V_{h}^{\text{V}}\in\mathbb{R}^{n\times d_{h}}$ denotes the key and value vectors of the corresponding skill in $V_{X}$, and $V_{h}^{\text{Q}}\in\mathbb{R}^{d_{h}}$ denotes the query projected from the set embedding. Using the dot-product attention mechanism [57], we calculate the similarity of each skill in each view as:
$V^{\text{E}}_{h}=V_{h}^{\text{Q}}(V_{h}^{\text{K}})^{T}/\sqrt{d_{h}}.$ (1)
Each entry $V^{\text{E}}_{h}\in\mathbb{R}^{N}$ models the interaction weight of the $i$-th skill with the $h$-th view.
While activation functions can be used to project interaction weights to the range [0, 1], the continuous interaction score hinders the network’s ability to model in an explicit binary manner. To address this issue, we introduce Gumbel-Sigmoid [58, 59] to enable differentiable skill selection. Specifically, we incorporate Gumbel noise with the Sigmoid function and output the selected skills for each view’s subset as:
$\begin{split}A^{\text{s}}_{h}=\frac{\exp((V^{\text{E}}_{h}+G_{0})/\tau)}{\exp((V^{\text{E}}_{h}+G_{0})/\tau)+\exp((G_{1})/\tau)},\\ u_{i}\sim\text{Uniform}(0,1),i\in\{0,1\},\\ G_{i}=-\log(-\log(u_{i})),i\in\{0,1\},\end{split}$ (2)
where $G_{0}$ and $G_{1}$ are two i.i.d Gumbel noises, the temperature $\tau$ controls the smoothness of the sampling function. $A^{\text{s}}_{h}$ becomes a multi-hot vector for selecting a subset $S_{h}$ when $\tau$ approaches 00 and becomes a continuous vector as $\tau\rightarrow\infty$. In the training loop, we gradually reduce $\tau$ to learn the multi-hot representation of the selected subset.
4.2.2 Skill-wise External Information Fusion